# Never-Ending Embodied Robot Learning

### v1

@@ 4.1 Implementation Details
Table 3: Comparison results on Living Room under the setting of 6-3. Red and Blue represents the highest results and runner-up.
Comparison Methods 1 2 3 4 5 6 7 8 9 10 11 12 All Imp.
@@
Buffer size 6-3 $(2~{}\mathrm{steps})$ 6-2 $(4~{}\mathrm{steps})$
1-6 7-12 All Avg. For. 1-6 7-12 All Avg. For.
$|\mathcal{M}|=2$ 42.0 41.3 41.7 44.9 4.4 32.0 30.7 31.3 37.2 18.8
$|\mathcal{M}|=4$ 44.9 42.2 43.6 47.6 7.7 45.8 35.3 40.6 43.5 10.9
$|\mathcal{M}|=6$ 50.0 36.7 43.3 45.0 2.2 50.0 34.7 42.3 45.4 8.0
Baselines. We conduct the comprehensive evaluation between our NBCagent and the following four methods: PerAct (Shridhar et al., 2023), GNFactor (Ze et al., 2023), Fine-Tuning and ER (Chaudhry et al., 2019). PerAct and GNFactor joint train all manipulation skills within one-stage training, for two datasets respectively, referred to as the upperbound. Fine-Tuning achieves continual learning by fine-tuning all parameters in perceiver model on novel skills. ER randomly stores old skill data to memory buff and replay them when detecting novel skills.
Training Details. In NERL, we assume that the agent undergoes initial learning through a set of manipulation skills, referred to as base task, while characterizing novel skills as incremental tasks. On Kitchen dataset, the base task includes 5 manipulation skills, and each incremental task consists of 1 manipulation skill (total 6 steps) and 5 manipulation skills (total 2 steps), marked as 5-1 and 5-5. Likewise, on Living Room dataset, 12 manipulation skills are divided into two NERL settings: 6-3 (total 3 steps) and 6-2 (total 4 steps). In addition, the LAMB (You et al., 2019) optimizer is applied for all methods with a initial learning rate of $5.0\times 10^{-4}$ and a batch size of 2. We utilize 100K training iterations for PerAct and GNFactor, 80K for base task training, 20K for incremental task training. We store a fixed 4 episodes of each old skill in $\mathcal{M}$ for ER and Ours. Additional training details are available in Appendix B.
Evaluation Metric. Following ERL methods (Ze et al., 2023; Goyal et al., 2023), we use the success score (%) $\mathcal{I}^{m}_{i}$ as the basic indicator for evaluation, where $\mathcal{I}^{m}_{i}$ represents the success score of $i$-th manipulation skill at $m$-th incremental task. Specifically, we first compute the mean success score after the last step for the base task (Base) $\mathcal{I}^{M}_{B}$, incremental tasks (Novel) $\mathcal{I}^{M}_{N}$ and all manipulation skills (All) $\mathcal{I}^{M}_{A}$. These metrics respectively reflect the robustness of old skill forgetting , the capacity of novel skill learning, as well as its overall performance. Additionally, we introduce a Avg. metric $A$ and For. metric $F$ to measure average performance and skill-wise forgetting rate over the whole NERL process, where $A=\frac{1}{M}\sum_{m=1}^{M}\mathcal{I}^{m}_{A}$ and $F=\frac{1}{N^{m}}\sum^{N^{m}}_{i=1}\underset{m\in\{1,...,M-1\}}{\mathrm{max}}(\mathcal{I}^{m}_{i}-\mathcal{I}^{M}_{i})$.
4.2 Comparison Performance
We present comparison results between our NBCagent and other methods on Kitchen and Living Room datasets in Tabs. 1, 2 and 3. As shown in Tab. 1, NBCagent significantly outperforms compared methods by $1.2\%\sim 51.2\%$ in terms of success score on base task and $5.1\%\sim 17.6\%$ on incremental tasks. This indicates that NBCagent can learn skill-specific and skill-shared knowledge, thereby effectively addressing old skill forgetting and novel skill learning. Furthermore, as presented in Tabs. 2 and 3, our model exhibits the highest mean success rate across all manipulation skills, improving by $1.2\%\sim 36.3\%$ and $3.6\%\sim 25.6\%$ respectively when compared to other methods. This demonstrates the effectiveness of our model in addressing the NERL problem. Additionally, our NBCagent achieves a large improvements about $2.5\%\sim 41.4\%$ and $0.1\%\sim 36.4\%$ in terms of Avg. and For. metrics, which suggests the robust and significant performance of NBCagent over the whole NERL process. Surprisingly, as shown in Tabs. 2 and 3, our NBCagent performs better than joint training methods, specifically PerAct and GNFactor, providing additional evidence to support the efficacy of our model. The comparison results under other settings can be found in Appendix C.
4.3 Ablation Studies
To evaluate effectiveness of each module in our NBCagent, we eliminate them one by one and present results in Tabs. 1, 2 and 3. Compared to Ours, the scores of Ours-w/oSRD on both base task and all tasks are dropped by $3.8\%\sim 12.5\%$ and $0.8\%\sim 5.8\%$ respectively. This indicates that SRD can effectively learn skill-shared knowledge to tackle old task forgetting. In addition, Ours-w/oSRD outperforms Ours-w/oSEP&SRD on incremental tasks by $1.3\%\sim 11.4\%$, which demonstrates that SEP benefits our model in learning novel skills by performing skill-specific knowledge learning. Furthermore, to evaluate the effectiveness of SSR module, we visualize the rendering results in Fig. 4. It suggests that our SSR module can efficiently complete skill-shared semantics under the supervision of novel view and diffusion features, thereby achieving an improvement about $0.5\%\sim 3.1\%$ in terms of Avg. compared with ER. The visualization results in Fig. 3 also shows the effectiveness of our model to tackle the NERL problem. We also explore the impact of various sizes of memory buffer $\mathcal{M}$. As shown in Tab. 4, with $|\mathcal{M}|=6$, the forgetting rate of our model notably reduced by $1.1\%\sim 10.8\%$ in comparison to $|\mathcal{M}|=4$ and $2$, respectively. This indicates that increasing the memory size significantly addresses catastrophic forgetting but also incurs a larger memory load.
@@ 5 Conclusion
In this paper, we explore a pioneering Never-ending Embodied Robot Learning (NERL) problem and propose a novel NBCagent to continually learn skill-wise knowledge. Specifically, we propose a skill-specific evolving planner to decouple the skill-wise knowledge to effectively learning novel skills. In addition, we design a skill-shared semantics rendering module and skill-shared representation distillation module to tackle catastrophic forgetting on old skills from semantics and representation aspects. We develop two NERL benchmarks and expensive experiments on them verify the effectiveness of our NBCagent against baselines.
Appendix A Dataset Description
@@
In order to emulate the operational environments of domestic robots, we design two NERL benchmark datasets using RLBench, called Kitchen and Living Room. In particular, we assemble Kitchen by gathering 10 skills that are pertinent to kitchen settings, and we construct Living Room, which includes 12 skills that are related to living room situations. Each skill includes a training set of 20 episodes and a test set of 25 episodes. We provide examples of brief overviews of these skills and the number of keyframes for each skill in Tab. 5. In addition, the Skill No. represents the corresponding skill order mentioned in the Experiment section. The types of skill variables include: the color, position, and size of the target object. The skill description may randomly sample from twenty colors: red, maroon, lime, green, blue, navy, yellow, cyan, magenta, silver, gray, orange, olive, purple, teal, azure, violet, rose, black, and white; and two sizes: short and tall. The position variable is skill-specific. For example, in the skill ‘place wine at rack location’, the target object ‘rack’ has three positions: ‘middle’, ‘left’, ‘right’.
@@
Skill No. Skill in Kitchen Language Example Avg. Keyframes
1 close microwave ”close microwave” 2.3
2 meat off grill ”take the steak off the grill” 6.0
3 open grill ”open the grill” 4.5
4 open wine bottle ”open wine bottle” 3.3
5 pick up cup ”pick up the red cup” 3.8
6 turn tap ”turn left tap” 3.0
7 place wine at rack location ”stack the wine bottle to the left of the rack” 6.1
8 put knife on chopping board ”put the knife on the chopping board” 5.1
9 stack wine ”stack wine bottle” 7.4
10 take plate off colored dish rack ”take plate off the red colored rack” 6.3
Skill No. Skill in Living Room Language Example Avg. Keyframes
1 close door ”close the door” 4.8
2 close laptop lid ”close laptop lid” 6.0
3 hang frame on hanger ”hang frame on hanger” 5.3
4 lamp on ”turn on the light” 3.5
5 open drawer ”open the bottom drawer” 4.8
6 open window ”open left window” 6.0
7 push buttons ”push the white button” 4.8
8 put item in drawer ”put the item in the top drawer” 14.3
9 put rubbish in bin ”put rubbish in bin” 5.0
10 sweep to dustpan of size ”sweep dirt to the tall dustpan” 6.2
11 take usb out of computer ”take usb out of computer” 3.3
12 water plants ”pour some water on the plant” 6.3
Table 6: Comparison results on Kitchen under the setting of 5-5. Red and Blue represents the highest results and runner-up.
@@
Table 7: Comparison results on Living Room under the setting of 6-2. Red and Blue represents the highest results and runner-up.
@@
Appendix B Training Details
As shown in Tab. 8, we give the hyper-parameters used in NBCagent. The overall learning rate of the training process is 0.0005, optimized by the LAMB algorithm. All methods employ the same methodology to construct a voxel grid of size $100^{3}$. During the NeRF training, we utilize an additional 19 camera perspectives to furnish supervisory information. For NERL, we conduct two NERL settings in each of the two scenarios, which are 5-1, 5-1 in Kitchen, and 6-3, 6-2 in Living Room. In SEP, NBCagent detects novel skills through the cosine similarity of the skill language description feature vectors. We set the judgment threshold $\delta$ as 0.8 for our experiments.
Table 8: Hyper-parameters used in NBCagent
Variable Name Value
image size 128 × 128 × 3
input voxel size 100 × 100 × 1000
batch size $B$ 2
optimizer LAMB
learning rate 0.0005
number of transformer blocks 6
number of sampled points for NERagent 64
number of latents in Perceiver Transformer 2046
dimension of Stable Diffusion features 512
dimension of CLIP language features 512
hidden dimension of NeRF blocks 512
size of LoRA layers $N^{r}$ 10
base task iterations 80K
incremental task iterations 20k
SSR loss weight $\lambda_{1}$ 0.1
SSR loss function $\mathcal{L}_{s}$ MSE-loss
SRD loss weight $\lambda_{2}$ 0.2
SRD loss function $\mathcal{L}_{srd}$ KL divergence
SRD temperature $\mathcal{T}$ 3
novel skill threshold $\delta$ 0.8
size of memory buffer $\mathcal{M}$ 4
number of base skill in living room 6
number of incremental skill in living room 3, 2
number of base skill in kitchen 5
number of incremental skill in kitchen 1, 5
Appendix C More Comparison Experiments
Tabs. 6 and 7 show the detailed scores of each model in two scenarios under the NERL settings of 5-5 and 6-2.With the configuration of memory buffer size $|\mathcal{M}|=4$, the average score of NBCagent across all skills still surpasses other models, approximately between $0.4\%\sim 23.8\%$ and $0.6\%\sim 27.9\%$. This indicates that NBCagent performs the best in solving NERL problems. Fig. 5 shows the visualization results of more skills among various models. We observe that some skills are not difficult to learn the general actions, but when the skill requires the target object to be in a specific location, the agent has difficulty completing it accurately; for example, ’open the drawer’ is not difficult, but when the requirement is ’open the top drawer’, the performance of NBCagent is better than other models.

### v2

@@ 4.3 Ablation Studies
Table 4: Comparison results on Living Room under the setting of 6-3.
Skill ID 1 2 3 4 5 6 7 8 9 10 11 12 All Imp.
@@ 5 Conclusion
In this paper, we explore a pioneering Never-ending Behavior-cloning Robot Learning (NBRL) problem and propose a novel NBAgent to continually learn skill-wise knowledge. Specifically, we design a skill-shared semantic rendering module and a skill-shared representation distillation module to tackle 3D scene representation overlooking on past encountered skills. Meanwhile, we propose a skill-specific evolving planner to decouple the skill-wise knowledge to effectively learning novel skills and tackling catastrophic forgetting. We develop two NBRL benchmarks to demonstrate the manipulation accuracy, which verifies the effectiveness of our NBAgent against baselines.
Appendix A Supplemental Material
This supplemental material introduces additional details not mentioned in the main paper. Overall, it mainly includes the following aspects:
(a) More Dataset details of Kitchen and Living Room in Sec. A.1.
(b) More implementation details of our proposed NBAgent in Sec. A.1.
(c) More comparison experiments in Sec. A.3.
(d) Analysis of limitations Sec. A.4.
(e) Analysis of societal impact in Sec. A.5.
Table 5: Some important notations and definitions used in this paper.
Notations Definitions Notations Definitions
$\mathcal{T}$ The skill data stream $\mathbf{Y}_{c}$ The input image with ground truth color
$\mathcal{T}^{m}$ The $m$-th continual skill learning task $\mathbf{Y}_{c,t}$ A noisy image
$\mathcal{D}^{m}_{i}$ The demonstration of $i$-th skill in $m$-th task $\mathbf{M}$ The 2D semantic map
$\mathbf{a}$ Action space $\Theta_{u}$ The Stable Diffusion model
$\mathbf{r}$ Structured observation $\mathcal{E}_{v}$ The VAE encoder for image $\mathbf{Y}_{c}$
$\mathbf{l}$ Language instruction $\mathbf{F}_{s}$ The rendered vision-language semantic feature
$\mathbf{Q}^{m}$ The Q-function head in $m$-th continual skill learning task $\hat{\mathbf{F}_{s}}$ The diffusion vision-language semantic feature
$\Theta^{m}_{p}$ The perceiver model in $m$-th continual skill learning task $t$ The diffusion process step
$\Theta^{m}_{n}$ The NeRF model in $m$-th continual skill learning task $\mathbf{l}_{p}$ The language prompt modified from $\mathbf{l}$
$\mathcal{F}_{\Theta^{m}_{n}}$ The neural rending function of $\Theta^{m}_{n}$ $\mathcal{E}_{c}$ The language encoder from CLIP
$\mathbf{v}$ The 3D voxel observation $\mathbf{l}_{s}$ A skill-specific language semantic embedding
$\mathbf{v}_{s}$ The 3D voxel feature $\mathcal{B}$ The adaptive language semantic bank
$\mathbf{x}$ The 3D input point $\mathcal{C}$ The cosine similarity matrix
$\sigma$ Differential density $\mathrm{nonzero}(\cdot)$ An operation employed to calculate the number of nonzero vectors in $\mathcal{B}$
$\mathbf{d}$ The unit viewing direction $h$ A skill-wise code
$\mathbf{r}$ Camera ray $\mathbf{S}$ Dynamic skill-specific latent space
$\mathcal{R}$ The set of all camera rays $\mathbf{F}_{c}$ The cross-attention feature
$\mathbf{o}$ Camera origin $\rho$ The softmax function
$\mathbf{c}$ The color of RGB $\mathbf{e}$ The encoded embeddings of input language instruction
$\mathbf{s}$ The semantic feature $\mathbf{W}$ The MLP projection layer in Transformer block
$\mathbf{C}$ A RGB image $\mathbf{k}^{m}_{j}$ The $j$-th keyframe in a skill, is same with $\mathbf{k}^{m,i}_{j}$
$\hat{\mathbf{C}}$ The pseudo ground truth $\mathbf{v}^{m}_{r,j}$ A keyframe on voxel space get through $\Theta^{m}_{p}$
$N^{d}$ The number of skills in a continual skill learning task $\mathbf{Y}^{m}_{j}$ The ground truth for prediction the next keyframe
$N^{b}$ The number of voctors in $\mathcal{B}$ $\hat{\mathbf{Y}^{m}_{j}}$ The soft label extracted by the perceiver model $\Theta^{m-1}_{p}$
$N^{s}$ The number of learned skills $\mathbf{X}$ The input feature for projection layers in Transformer
$N^{l}$ The learnable latent vector quantity $D$ The dimension of latent vectors in projection layers in Transformer
$N^{p}$ The number of encoded embeddings of input data $D^{s}$ The dimension of $\mathbf{l}_{s}$
$N^{e}$ The number of encoded embeddings of the input language instruction
$N^{r}$ The size of LoRA layers
A.1 Dataset Description
@@
In order to emulate the operational environments of domestic robots, we design two NBRL benchmark datasets using RLBench, called Kitchen and Living Room. In particular, we assemble Kitchen by gathering 10 skills that are pertinent to kitchen settings, and we construct Living Room, which includes 12 skills that are related to living room situations. Each skill includes a training set of 20 episodes and a test set of 25 episodes. We provide examples of brief overviews of these skills and the number of key frames for each skill in Tab. 6 and visualization in Figs. 4-5. In addition, the No. represents the corresponding skill order mentioned in the Experiment section. The Avg. Kf describes the average number of key frames for the skill over 20 episodes in training set.
@@
No. Skill in Kitchen Language Example Avg. Kf
1 close microwave "close microwave" 2.3
2 meat off grill "take the steak off the grill" 6.0
3 open grill "open the grill" 4.5
4 open wine bottle "open wine bottle" 3.3
5 pick up cup "pick up the red cup" 3.8
6 turn tap "turn left tap" 3.0
7 place wine "stack the wine bottle to the left of the rack" 6.1
8 put knife on "put the knife on the chopping board" 5.1
9 stack wine "stack wine bottle" 7.4
10 take plate off "take plate off the red colored rack" 6.3
No. Skill in Living Room Language Example Avg. Kf
1 close door "close the door" 4.8
2 close laptop lid "close laptop lid" 6.0
3 hang frame "hang frame on hanger" 5.3
4 lamp on "turn on the light" 3.5
5 open drawer "open the bottom drawer" 4.8
6 open window "open left window" 6.0
7 push buttons "push the white button" 4.8
8 put item in drawer "put the item in the top drawer" 14.3
9 put rubbish in bin "put rubbish in bin" 5.0
10 sweep to dustpan "sweep dirt to the tall dustpan" 6.2
11 take usb out "take usb out of computer" 3.3
12 water plants "pour some water on the plant" 6.3
Figure 4: Visualization of manipulation skills in dataset Kitchen.
Figure 5: Visualization of manipulation skills in dataset Living Room.
Table 7: Hyper-parameters used in NBAgent
Variable Name Value
image size 128 × 128 × 3
input voxel size 100 × 100 × 100
batch size $B$ 2
optimizer LAMB
learning rate 0.0005
number of transformer blocks 6
number of sampled points for NBAgent 64
number of latents in Perceiver Transformer 2046
dimension of Stable Diffusion features 512
dimension of CLIP language features 512
hidden dimension of NeRF blocks 512
size of LoRA layers $N^{r}$ 10
base task iterations 80K
incremental task iterations 20K
SSR loss weight $\lambda_{1}$ 0.1
SSR loss function $\mathcal{L}_{SSR}$ MSE-loss
SRD loss weight $\lambda_{2}$ 0.2
SRD loss function $\mathcal{L}_{SRD}$ KL divergence
SRD temperature $\mathcal{T}$ 3
novel skill threshold $\delta$ 0.8
size of memory buffer $\mathcal{M}$ 4
number of base skills in living room 6
number of novel skills in living room 3, 2
number of base skills in kitchen 5
number of novel skills in kitchen 1, 5
Table 8: Comparison results on Kitchen under the setting of 5-5.
@@
Table 9: Comparison results on Living Room under the setting of 6-2.
@@
A.2 Implementation Details
As shown in Tab. 7, we give the hyper-parameters used in NBAgent. One agent is trained with two NVIDIA A40 GPU for 2 days (80K + 40K iterations under the 6-3 setting in Living Room) with a batch size of 2. The overall learning rate of the training process is 0.0005, optimized by the LAMB algorithm. All methods employ the same methodology to construct a voxel grid of size $100^{3}$. During the NeRF training, we utilize an additional 19 camera perspectives to furnish supervisory information. For NBRL, we conduct two NBRL settings in each of the two scenarios, which are 5-5, 5-1 in Kitchen, and 6-3, 6-2 in Living Room.
A.3 More Comparison Experiments
Tabs. 8 and 9 show the detailed scores of each model in two scenarios under the NBRL settings of 5-5 and 6-2.With the configuration of memory buffer size $|\mathcal{M}|=4$, the average score of NBAgent across all skills still surpasses other models, approximately between $0.4\%\sim 23.8\%$ and $0.6\%\sim 27.9\%$. This indicates that NBAgent performs the best in solving NBRL problems. Fig. 6 shows the visualization results of more skills among various models. We observe that some skills are not difficult to learn the general actions, but when the skill requires the target object to be in a specific location, the agent has difficulty completing it accurately; for example, ’open the drawer’ is not difficult, but when the requirement is ’open the top drawer’, the performance of NBAgent is better than other models.
@@ A.3 More Comparison Experiments
A.4 Limitation
While we have found that NBAgent achieves state-of-the-art results on the Never-ending Behavior-cloning Robot Learning (NBRL) problem, we have also identified the following limitations: 1. NBAgent collects keyframes as learning targets using predefined rules in the demonstration dataset, but these rules are not universally applicable to all manipulation tasks. 2.NBAgent cannot continuously learn skills that require continuous actions as it predicts actions based on time-discrete control.
A.5 Social Impact
In practical multi-task robot learning, we explore the Never-ending Behavior-cloning Robot Learning (NBRL) problem and propose a novel NBAgent. The goal is to continually learn new skills in an incremental manner, allowing for unique customized operational abilities. For example, robots can learn new skills by imitating specific demonstration datasets tailored to the owner’s needs. Additionally, we address catastrophic forgetting in continual skill learning by transferring skill-shared semantic of 3D scenes using a skill-shared semantic rendering module and a skill-shared representation distillation module.
Typically, robots need strong learning capabilities to adapt to various work scenarios effectively. Having the ability for continual learning allows robots to provide better user experiences. Furthermore, this approach enables robots to learn specific operational skills based on the diverse needs of different user groups. For instance, household robots need to learn operational tasks in various contexts. Therefore, the proposed Never-ending Behavior-cloning Robot Learning problem is worth investigating.